After the driver passed by the school, he changed his e. speed again. The graph of speed versus time was plotted again i. Is the speed increasing or decreasing? Why did the driver change the speed? What is the type of motion in this case Time 5. A car takes 1mn 2s to cover a road 0.8 km long. The maximum allowed a. Did the driver violate the traffic regulations? b. How long does it take him to cover the road at 48km/h 6. A train 300m long passes through a tunnel 210m long. The interval between the entrance of the locomotive and the exit of the last wagon is 17s. Calculate the average velocity of the train 7. Complete the following table: Moving object Average speed Average Distance Time interval speed (Km/h) 16mm 100m Snail Runner Car Airplane Maglev 4s 10s 30 min 2h 30min 50km/h 1800km 200km 400km/h 8. Finally, Ali and his family got a vacation. Theyve decided to go to the village. They woke up early in the morning at 5:00 AM. Theypackaged all their stuffs. The father drove the car at 5:45 AM. At 7:00 they reached Tyre. They had breakfast in the grandparents house. At 8:55 they headed to the village. At 9:50 they reached the village. At the end of
Answer #1

5)Given,

t = 1 min 2 s = 62 s ; d = 0.8 km = 800 m ; v = 50 km/h = 13.89 m/s

a)We know that

v(driver) = d/t = 800/62 = 12.9 m/s

No he did not violate the traffic regulations as v(driver) = 12.9 m/s < v (13.89 m/s)

b)we know that

v = d/t => t = d/v

v = 48 km/h = 13.33 m/s

t = 800/13.33 = 60.02 s

Hence, t = 60.02 s